OpenStudy (anonymous):

The velocity function is v(t)= -t^2+5t-6 for a particle moving along a line. Find the distance traveled.

OpenStudy (anonymous):

I already got the displacement to be -67.5 I just need the distance traveled

OpenStudy (shayaan_mustafa):

differentiate with respect to t

OpenStudy (anonymous):

can you work it out?

OpenStudy (shayaan_mustafa):

\[\frac{dv}{dt}=\frac{d(-2t ^{2})}{dt}+5\frac{d(t)}{dt}-\frac{d(6)}{dt}\]\[x=-4t+5\]it your required.

OpenStudy (anonymous):

so now what

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Can someone please help me?

OpenStudy (experimentx):

Distance travelled = \(\large \int_{t_1}^{t_2}V(t)dt\)

OpenStudy (shayaan_mustafa):

damn.. i did opposite.. @anthonys it is acceleration. yes yes.. you must integrate to find distance.

OpenStudy (anonymous):

first put v(t)=0 you will get the time when it stops then integrate v(t) and put the value of t you will get your answer

OpenStudy (shayaan_mustafa):

@experimentX thnx for correction.

OpenStudy (experimentx):

happens sometime ... happens to me all the time.